OpenStudy (anonymous):

What is the y-intercept of the equation? y = 2.1(0.15)^x a.) 0.15 b.) 1 c.) 2.1

Directrix (directrix):

c.) 2.1

Directrix (directrix):

To get the y-intercept, let x be 0. y = 2.1 (0.15)^x y = 2.1 * .15^0 y = 2.1 * 1 y = 2.1